Pilot 2.0 (Feb 27) Migration Report

posted Feb 29, 2012, 9:42 AM by Rita Girardi

We will provide daily updates after each major phase of the project until technical issues and support issues return to a moderate level.

Pilot 2.0 occurred this past weekend. Participants in this migration included: ITS, the Google Guides for units migrating in May and designated LSA staff.

Pilot Participants

Users migrated: 821
Resources migrated: 61 ("Resources" include conference rooms)

Migration Summary

Since large calendar migration can take a long time to complete, we started the migration of 25 people with large calendars at 5:00 pm on Friday. Google migration activities for the rest of the staff began at 8:00 pm.

  • Messages succeeded/attempted: 12,762,596 / 12,770,371
  • Accuracy rate: 99.94%
  • Calendar events succeeded/attempted: 896,174 / 898,787
  • Accuracy rate: 99.71%

Support Summary

Transition Desk

  • Saturday, February 25 ... 9 tickets (mostly general usage)
  • Sunday, February 26 ... 6 tickets (mostly general usage)
  • Monday, February 27 ... 79 tickets (migration, resources, mobile, and general usage)

Google Guides

Student Google Guides and Project Team members were on-site at Administrative Services Building, Boyer, Arbor Lakes, Wolverine Tower, Hatcher Library, and Duderstadt Center. A total of 92 staff were provided assistance. The topics of the support questions that were covered with these people were:
  • Email: 100
  • Calendar: 47
  • Mobile device set-up: 23
  • Collaboration tools: 6

Issues

  • Some conference rooms had so many calendar items that they would not migrate. We had to remove old events from these calendars (e.g. removed the events before January 1, 2011).
  • Ownership of some resources did not migrate correctly. We are manually correcting these and fixed the script to avoid problems in future migrations.
  • Four people were on the list to be migrated but did not migrate. One person was hired after the Exchange list was generated. Their accounts were migrated on Monday.
  • Several IMAP mail migrations did not complete until Saturday night / Sunday morning due to the large number of messages. Affected users were contacted making them aware of the situation.
  • Exchange accounts were not set properly setup for approximately 50 users. Script had to be run to correct these accounts.
  • There were bounced emails for some users. This was a problem we thought we corrected after seeing it in a previous pilot. We need to revisit the mails that bounced to see what special situations we missed in our process.
  • Four users reported that voice mail did not go into email box. We're still investigating, but may be related to Google SPAM filtering.

Conclusion

Since the are no major open issues, we won't send any daily bulletins for Pilot 2.0. The next daily bulletin will be sent Monday, March 5 after the general availability of Google to faculty, staff and sutdents.
Comments